UTC 2011 May 16

Observer(s): T. Sargent, P. Grenz, A. Rakich, O. Kuhn, J. Hill (Tucson)
Telescope Operator: D. Huerta (LBTO)
Telescope Support:: M. Midkiff (LBTO)
Instrument Support: D. J. Morris (LBTO)

Summary

Preparations

Norm, Dan and Geno were running BP11 tests during the day.

Details

TCS Testing with BP11 (first half)

Lost first hour to wind.

Tom and Paul debugged problem(s) in MCS where on-source flag was always true. This took a long time.

John and Paul did the test of the M3 UMAC offsets. It took a long time to sort out because the sign of the offset was reversed, and we uncovered a couple other bugs in OSS. This took a long time. The other bugs were fixed the next day.

We did the Michele tests of the back calculation of coordinates in PCS.

Ilya could send IRTC presets on right, but had problems with LUCI presets on left.

We did just a few presets in monocular mode where we had stars.

SX Primary Mirror temperatures have reference junction problems.

Range Balancing Testing (second half)

Closed for wind at 7:05 UT, so no sky data.

John and Olga debugged the part of binorb.cl that sends commands to the telescope.

10:25 Give up for the night.

-- JohnHill - 16 May 2011
Topic revision: r3 - 18 May 2011, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback